What is the pH of a 0.1M solution of benzoic acid? ( Write answer to the hundredths place) KA(benzoic acid) = 6.46x10-5
Answer #1

Lets see the dissociation reaction of benzoic acid

C6H5COOH \rightleftharpoons H+ + C6H5COO-  

we have

[C6H5COOH ] = 0.1 M

Ka = 6.46 \times 10-5  

Lets see the ICE table for above reaction

..........................C6H5COOH....\rightleftharpoons.....H+ ......+ ....C6H5COO-  

Initial.....................0.1 M.....................0......................0

Change ................-x...........................+x....................+x

At Equilibrium.....0.1-x........................x......................x

The equilibrium constant for above reaction is

Ka = [H+][C6H5COO-] / [C6H5COOH]

Hence

6.46 \times 10-5 = (x)(x) /(0.1-x) = x2 / (0.1-x)

as x is too small than 0.1 then 0.1-x = 0.1

so

x2 / 0.1 = 6.46 \times 10-5

x2 = 0.1 \times 6.46 \times 10-5

x2   = 6.46 \times 10-6

by taking square roots on both the side , we get

x = 0.0025

then

[H+] = x = 0.0025 M

now we know that

pH = -log[H+]

= -log(0.0025)

= 2.60

so our answer is

pH = 2.60